OpenStudy (republican31):

2+3x=4y+5z ______ 6 for Z... All answers are some variation of z=12+18x+24y _____________ 5 I'm doing the line system because I was afraid it could be confused with just 5z/6 instead of 4y+5z all over 6.

jimthompson5910 (jim_thompson5910):

you can type it like this 2+3x=(4y+5z)/6

OpenStudy (republican31):

Okay, I just wasn't sure if putting those in ()'s changed the equation at all.

jimthompson5910 (jim_thompson5910):

anyways, first you multiply both sides by 6 2+3x=(4y+5z)/6 6*(2+3x)=6*(4y+5z)/6 12+18x=4y+5z

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (republican31):

12+18x=24y+30z?

OpenStudy (republican31):

Then 12+18x-24y=30z, right?

OpenStudy (republican31):

Should I not have multiplied 5 and 6? Or what am I supposed to do now?

jimthompson5910 (jim_thompson5910):

when you multiply (4y+5z)/6 by 6, the two '6' terms divide to 1 1 times anything is the same number, so the '6's go away more or less

jimthompson5910 (jim_thompson5910):

(4y+5z)/6 will turn into 4y+5z after you multiply the right side by 6

OpenStudy (republican31):

Okay. So z=(12+18x-4y)/5.

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (republican31):

Thank you!

jimthompson5910 (jim_thompson5910):

sure thing